OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 3y = 7 3x - 5y = -23 -1 1 -4 4

OpenStudy (anonymous):

eliminate x

OpenStudy (anonymous):

I got stuck at 3y=-5x+7

OpenStudy (anonymous):

oh you doing sub method :)

OpenStudy (anonymous):

it would be easier to use addition/subtraction of equations

OpenStudy (anonymous):

15x+9y = 21 15x - 25y = -115 subtract 34y = 136 y = 136/34 y = 4 plug 4 back into original equations 5x +12 = 7 5x = -5 x = -1

OpenStudy (anonymous):

okay thank you! :)

OpenStudy (anonymous):

15x+9y = 21 15x - 25y = -115 subtract 34y = 136 y = 136/34 y = 4 stop there :P